OpenStudy (emsxx3_):

Use △DEF, shown below, to answer the question that follows: What is the value of x rounded to the nearest hundredth? Type the numeric answer only in the box below.

OpenStudy (emsxx3_):

@agent0smith

OpenStudy (emsxx3_):

please help need answer quick

OpenStudy (agent0smith):

\[\large \cos 49 = \frac{ x }{ 55 }\]

OpenStudy (emsxx3_):

i did it on the calculator and got 41.5085.

OpenStudy (emsxx3_):

I used an online scientific calculator. SOH so X/55. 49 sin = .754 then I multiplied by 55 and got 41.5085.

OpenStudy (emsxx3_):

is it correct?

OpenStudy (agent0smith):

See my post above.

OpenStudy (emsxx3_):

okay whats ?

OpenStudy (emsxx3_):

whats X ?

OpenStudy (agent0smith):

x is what you're finding. Looks like you did it almost right, except you used sin not cos.

OpenStudy (emsxx3_):

how do you find x with cos

OpenStudy (agent0smith):

55 cos 49 = x